• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ole Hansen

    Log entry time 07:48:33 on June 9,2000

    Entry number 43211

    keyword=Shift Summary

    Shift started waiting for beam.

    1:05 ... Beam ready
    1:20 ... 15cm LH2 Target in place, 5uA for spot check, ok
    1:25 ... asking for 50uA, accelerator has some trouble getting there.
    1:45 ... Start test run #1823 w/50uA, somewhat unstable, E-arm momentum
    ............still ramping to p=3.957 GeV.
    1:48 ... Q3 Electron overtemperature trip! Does not clear.
    1:58 ... Start test run #1824. E-arm still down, so we prescale it. Mark Jones
    ........... is mostly interested in H-arm (which is set for electrons).
    2:15 ... End run #1824. Replay for H-arm checks out fine. Moving in H-arm
    ........... sieve slit. FF on.
    2:19 ... Start run #1825 - optics data for H-arm
    3:30 ... Asking for 80uA. We get it.
    3:38 ... Changing to 12C target because we don't see the sieve slit with LH2.
    3:49 ... Start run #1846 - 12C, 50uA, no raster, H-arm only
    3:57 ... End run - looks like unusable data too.
    4:02 ... Starting y_tgt studies. Run 1847, 12C, 6msr collimators, H-arm only.
    4:09 ... Moving in 4cm dummy
    4:17 ... Start run # 1848, 5uA, 4cm dummy
    4:27 ... Beam goes away due to injector problems.
    5:10 ... While we wait for beam, we take a short access to reset Q3 on the E-arm.
    6:30 ... MCC calls - still clueless about the cause of the injector problem.
    ........... We decide to ramp down the hadron spectrometer in preparation for
    ........... the planned polarity change.
    7:40 ... MCC calls - a klystron blew - they're going to replace it, will be
    ........... at least 1 hour.